OpenStudy (anonymous):

integrate 2x/sqrt(1-x^2) dx from 0 to 1/2

myininaya (myininaya):

try a substitution \[\text{let } u=1-x^2\]

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

kids ok? :)

OpenStudy (anonymous):

haha yes, they are

OpenStudy (anonymous):

im stuck.

OpenStudy (anonymous):

i have \[-1/2\int\limits_{1}^{3/4}du/\sqrt{u} du\]

OpenStudy (anonymous):

i dont think the top of the function is right, it was the value i had for du but idk.....

OpenStudy (anonymous):

I would do trig substitution here.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

\[ \int\frac{2x}{\sqrt{1-x^2}}dx\\ x=\sin u\\ dx=\cos u\ du\\ \int\frac{2\sin u\cos u}{\sqrt{1-\sin^2u}}du=\int\frac{2\sin u\cos u}{\sqrt{\cos^2u}}du=\int\frac{2\sin u\cos u}{\cos u}du=\int2\sin u\ du \]

OpenStudy (anonymous):

how would you know to use that, since there is no trig functions in the equation.

OpenStudy (anonymous):

Any equation with a denominator like that is often a good candidate for a trig substitution. You'll be able to recognize them after enough practice.

OpenStudy (anonymous):

it looks like inverse sin

OpenStudy (anonymous):

@nbouscal it's very simple integral She just forgot to take root of the upper limit, it should be sqrt (3) /2

myininaya (myininaya):

much easier to do u=1-x^2

OpenStudy (anonymous):

Eh, there are plenty of methods to do any given integral. I find trig substitution easiest because I can usually do it in my head.

OpenStudy (anonymous):

the derivaitve if available so i second that motion^

myininaya (myininaya):

\[u=1-x^2 => du=-2x dx => -du =2 x dx\] \[\int\limits_{1-0^2}^{1-(\frac{1}{2})^2} \frac{-du}{\sqrt{u}}\] \[\int\limits_{1}^{\frac{4}{4}-\frac{1}{4}}-u^\frac{-1}{2} du\] \[-\int\limits_{1}^{\frac{3}{4}}u^{-\frac{1}{2}} du\]

OpenStudy (anonymous):

Yeah, I guess it's pretty easy that way too. Meh. Either way works fine.

myininaya (myininaya):

I wasn't saying your way didn't work. I was just saying the way I chose is much easier

OpenStudy (anonymous):

Matter of opinion.

OpenStudy (anonymous):

would there be a -1/2 in front of the integral?

myininaya (myininaya):

lol ok. You learn trig sub after algebraic sub But okay I guess it is sorta an opinion thing

myininaya (myininaya):

@jackiefml what does du equal? du=-2x dx so -du=2x dx Isn't 2x dx on top So isn't that -du?

OpenStudy (anonymous):

yes, but normally i have to separate the coeffecient from the dx.

myininaya (myininaya):

ok let me know if you run into anymore trouble

OpenStudy (anonymous):

i integrate next to get rid of the du?

myininaya (myininaya):

 This symbol in the picture means you need to integrate with respect to u yes

OpenStudy (anonymous):

ok so id get.... \[2\sqrt{u}\]

myininaya (myininaya):

don't forget the negative

OpenStudy (anonymous):

ok

myininaya (myininaya):

Then evaluate your antiderivative at the upper limit - evaluate it at the lower limit

myininaya (myininaya):

and if you prefer @nbouscal 's way then look at the attachment I did it both ways